# Emulating Self-attention with Convolution for Efficient Image Super-Resolution

### v1

@@ Abstract
In this paper, we tackle the high computational cost of transformers for lightweight image super-resolution (SR). Motivated by the observations of self-attention’s inter-layer repetition, we introduce a convolutionized self-attention module named Convolutional Attention (ConvAttn) that emulates self-attention’s long-range modeling capability and instance-dependent weighting with a single shared large kernel and dynamic kernels. By utilizing the ConvAttn module, we significantly reduce the reliance on self-attention and its involved memory-bound operations while maintaining the representational capability of transformers. Furthermore, we overcome the challenge of integrating flash attention into the lightweight SR regime, effectively mitigating self-attention’s inherent memory bottleneck. We scale up window size to 32$\times$32 with flash attention rather than proposing an intricated self-attention module, significantly improving PSNR by 0.31dB on Urban100$\times$2 while reducing latency and memory usage by 16$\times$ and 12.2$\times$. Building on these approaches, our proposed network, termed Emulating Self-attention with Convolution (ESC), notably improves PSNR by 0.27 dB on Urban100$\times$4 compared to HiT-SRF, reducing the latency and memory usage by 3.7$\times$ and 6.2$\times$, respectively. Extensive experiments demonstrate that our ESC maintains the ability for long-range modeling, data scalability, and the representational power of transformers despite most self-attentions being replaced by the ConvAttn module.
@@ 1 Introduction
Therefore, transformers have gained significant attention in the SR task since they have demonstrated superior performance over Convolutional Neural Networks (CNNs) while requiring lower FLOPs and fewer parameters. By capturing long-range dependencies and performing input-dependent weighting through self-attention, transformers exhibit high representational capacity that enhances performance, especially as the training data volume increases [11, 3]. Accordingly, numerous transformer-based methods have been proposed [55, 45, 37] to capitalize on the advantages of self-attention while reducing FLOPs and parameter counts.
Figure 1: Comparison of performance, latency, and memory usage on the Urban100$\times$4 dataset. Our methods are evaluated against representative SR models, including CNNs, transformers, and SSMs.
@@
In this paper, we aim to design an efficiency-driven transformer tailored for lightweight SR tasks, achieving both improved performance and memory overhead. We start by conducting a preliminary analysis of self-attention, observing that the similarity modeling performed by self-attention and resulting extracted features remain highly consistent across several layers, as shown in Figure 2. This finding exhibits that self-attention may extract overlapping features, suggesting the possibility of reducing computational overhead without compromising representational capability by utilizing efficient alternatives. Building on this finding, we establish a design strategy that retains self-attention only in the first layer of each block while replacing the remaining layers with our proposed efficient alternative, the Convolutional Attention (ConvAttn) module. To effectively emulate self-attention’s long-range modeling and instance-dependent weighting, ConvAttn operates with a twofold mechanism. First, it simplifies self-attention’s long-range interaction by applying convolution with a shared 13$\times$13 large kernel throughout the entire network, targeting only a subset of channels. Second, dynamic kernels are generated to capture input-dependent weighting, mimicking the adaptive nature of self-attention. By combining these components, ConvAttn significantly reduces the reliance on memory-intensive self-attention while maintaining the representational capability of transformers.
With most self-attention layers replaced by ConvAttn, we leverage this efficiency to further enhance the remaining self-attention layers. Specifically, we enlarge the window size of self-attention, significantly improving performance with only a slight increase in FLOPs. However, increasing the window size leads to an enlarged score matrix, substantially raising peak memory usage. To address this, we introduce flash attention [7, 10] into lightweight SR tasks, to avoid materializing the score matrix. Our optimized implementation allows us to scale up the window size to 32$\times$32 while reducing latency and memory usage by 16$\times$ and 12.2$\times$, respectively, as illustrated in Table 1.
@@
  • We mitigate the memory overhead of self-attention by integrating flash attention, thereby enlarging window size without excessive memory usage. To the best of our knowledge, this is the first successful application of flash attention in lightweight SR tasks.
@@ 2 Related Work
Traditional CNNs Early deep learning‐based SR research [9, 22, 29, 58, 57] primarily employed convolution that extracts local features from images. These approaches typically relied on stacked convolutions, which often resulted in over‐parameterization. To address this issue, numerous studies [21, 42, 27, 36] have shared their weights across multiple layers to reduce parameter size. However, many of them suffer from performance degradation since the input-independent weights are shared across multiple layers. In this paper, we share a large kernel through the entire network while leveraging input-dependently generated dynamic DWC [18] by each layer in parallel. This approach efficiently mitigates the parameter growth of the large kernel while maintaining each layer’s representational power.
@@
Transformers Pioneering studies that significantly reduce computational costs by computing self-attention within local windows [28] or channel-wise [51] have sparked considerable interest in applying transformers [44, 11] to the SR task. Subsequent studies have proposed methods that enhance the self‐attention mechanism of transformers to strengthen their long‐range modeling capabilities while reducing both FLOPs and parameters [55, 45, 5, 62, 54, 56, 37]. Nonetheless, transformers remain challenging to deploy due to the excessive memory access caused by self-attention and involved memory-bound operations. In this paper, we address this limitation by replacing self‐attention with convolutions that emulate its advantages, thereby preserving the advantages of transformers while reducing reliance on self-attentions and associated memory-bound operations.
@@
Large Kernel CNNs and State-Space Models Several recent studies have highlighted the computational burden of transformers and investigated alternative core operators to replicate their advantages. Specifically, CNN-based models [41, 61, 49] attempt to mimic the advantages of transformers by leveraging large kernels and pixel attention [59], while State-Space Models (SSM) [16, 17] adapt their verified long-range modeling abilities from 1D sequence domain [14, 13] to the 2D image domain. However, these approaches often overlook either the exceptional representational capability brought by self-attention’s sophisticated similarity modeling or the additional memory access caused by supplementary mechanisms to adapting 2D domains like multi-directional or attentive scanning. In this paper, we fully leverage the advantages of both self-attention and convolution while utilizing flash attention [7, 10] to address the memory access of self-attention without relying on complex additional mechanisms.
@@
Figure 2: Visualized inter-layer similarities of the feature extracted by self-attention ($F$) before added to skip path and corresponding attention map ($A^{avg}$). Both CKA [24] and cosine similarities are measured on the Urban100$\times$4 dataset. The window and head dimensions of the attention map are averaged for visualization.
@@ 3.1 Preliminary analysis
Transformers often achieve improved performance by utilizing self-attention’s long-range modeling capabilities and instance-dependent weighting. However, each self-attention incurs excessive memory overhead, resulting in high latency and memory usage. Interestingly, our preliminary research demonstrates that the sophisticated similarity modeling of self-attention ($A^{avg}$) and the resulting extracted features ($F$) exhibit a high degree of inter-layer similarities, with averages of 89% and 87%. This observation not only aligns with recent research [60] but also explains the success of studies that reduced computational cost by sharing the attention map across layers [55, 37]. Inspired by this observation, we investigate improving efficiency without degrading the representational capability by replacing self-attention with efficient alternatives.
@@ 3.2 Overall Structure
Our network is composed of four main components, as shown in Figure 3. First, the LR input image $I_{LR}\in\mathbb{R}^{H\times W\times 3}$ is transformed into a shallow feature $F_{0}\in\mathbb{R}^{H\times W\times C}$ through a 3$\times$3 convolution ($\mathrm{Conv}_{3\times 3}$). Here, $H$ and $W$ denote height and width $I_{LR}$, and $C$ denotes feature dimension. Next, $F_{0}$ and a large kernel $LK\in\mathbb{R}^{13\times 13\times 16\times 16}$ that is shared throughout the entire model are fed into the deep feature extractor $H$, and the resulting deep feature is added back to $F_{0}$ to produce the final feature $F\in\mathbb{R}^{H\times W\times C}$. In parallel, the image-wise skip module $S$ takes $I_{LR}$ as input to generate a skipped feature $F_{s}$, and the upsampler $U$ utilize both $F$ and $F_{s}$ as inputs to return the upsampled image $I_{SR}\in\mathbb{R}^{rH\times rW\times 3}$ where $r$ denotes upscaling factor. The entire process is formulated as follows:
@@ 3.3 ESC Block
Firstly, the $F_{i-1}$ is converted to the $F^{in}_{i}$ by layer normalization ($\mathrm{LN}$) and convolutional FFN ($\mathrm{ConvFFN}$[62] with kernel size three. Then, $\mathrm{LN}$ and self-attention ($\mathrm{SelfAttn}$) with a window size of 32 is employed to residual connection to extract $F_{i,0}$. Despite the large window size, we effectively reduce memory overhead by utilizing flash attention. Furthermore, since ConvFFN is placed before, the self-attention extract features considering local information without the complicated $QKV$ projection module. After establishing $F_{i,0}$, the block iterates residual paths composed of $\mathrm{ConvFFN}$ and proposed ConvAttn module ($\mathrm{ConvAttn}$) $M$ times. Although our methods do not utilize shifted-window self-attention, these convolution-based modules, after the self-attention, will extract inter-window features. Finally, the result of the last iteration $F_{i,M}$ go through $\mathrm{LN}$ and $\mathrm{Conv}_{3\times 3}$ and is added to $F_{i-1}$, resulting in $F_{i}$.
@@
Convolutional Attention Module The $\mathrm{ConvAttn}$ is designed to emulate self-attention’s two key advantages: capturing long-range dependencies and instance-dependent weighting. For this purpose, $\mathrm{ConvAttn}$ leverages to types of convolution formulated as:
@@
$F^{res}_{i,j}$ $=(F^{att}_{i,j}\circledast DK_{j})+(F^{att}_{i,j}\circledast LK),$
$F^{fuse}_{i,j}$ $=\mathrm{Conv}^{fuse}_{1\times 1}\bigl{(}\mathrm{Concat}(F^{fuse}_{i,j},F^{idt}_{i,j})\bigr{)},$
where $F^{\mathrm{CF}}_{i,j}$ denotes the features extracted by $\mathrm{ConvFFN}_{i,j}$ before. We first split the $F^{\mathrm{CF}}_{i,j}$ channel-wise, resulting in $F^{att}_{i,j}\in\mathbb{R}^{H\times W\times 16}$, which consists of the 16 channels in front, and $F^{idt}_{i,j}\in\mathbb{R}^{H\times W\times C-16}$, which contains the remaining part. Both large kernel and dynamic convolutions are only performed within $F^{att}_{i,j}$ to reduce memory access cost, unlike the methods that perform large kernel convolution to full channels [41, 49]. Next, dynamic depth-wise kernel $DK_{i,j}\in\mathbb{R}^{3\times 3\times 1\times 16}$ [18] is created by two layers of $\mathrm{Conv}_{1\times 1}$ with global average pooling ($\mathrm{GAP}$) and gaussian error linear unit activation ($\phi$). Since enlarging $DK$’s kernel size increases parameter size multiplied by the hidden dimension of the kernel estimator, we fix its kernel size to 3$\times$3. Then both $DK_{i,j}$ and $LK$ are convolved on $F^{att}_{i,j}$ and sum up those, resulting convolved feature $F^{res}_{i,j}$. Inspired by previous studies [55, 37] that shared attention maps across multiple layers and our preliminary analysis, we hypothesize that long-range interactions may not significantly change with varying layers. Therefore, we share $LK$ throughout the entire model, mitigating parameter growth and corresponding optimization difficulty [30, 4]. Consequently, harmonizing both $LK$ and $DK_{i,j}$ allows us to emulate self-attention’s advantages of capturing long-range dependencies and performing instance-dependent weighting with minimal increase in latency and parameter size. Finally, $F^{res}_{i,j}$ and $F^{idt}_{i,j}$ are concatenated channel-wise, followed by the last $\mathrm{Conv}_{1\times 1}$ to fuse them, resulting in fused feature $F^{fuse}_{i,j}$. In this way, both the local features extracted by $\mathrm{ConvFFN}$ and the long-range structural features extracted by $LK$ and $DK$ can be leveraged to capture multi-scale features.
While $DK_{i,j}$ can be structurally re-parameterized into $LK$ by padding both spatial and channel-wise to reduce FLOPs, we do not because it slows down latency. On the other hand, in situations where reducing FLOPs and parameter size is crucial rather than reducing latency, $LK$ can be decomposed depth-wise separable manner [19] into point-wise large kernel filter $LK^{\mathrm{c}}\in\mathbb{R}^{1\times 1\times 16\times 16}$ and depth-wise large kernel filter $LK^{\mathrm{s}}\in\mathbb{R}^{13\times 13\times 1\times 16}$. This decomposition redefines line 3 in Equation 4 as follows:
@@
In summary, the ConvAttn module efficiently emulates self-attention’s two advantages by leveraging a shared $LK$ and layer-specific $DK$s. By replacing self-attention and associated memory-bound operations with the proposed ConvAttn module, we significantly improve efficiency while maintaining the benefits of transformers.
Table 1: Comparison of latency and memory usage between naive self-attention and flash attention (FA) [7, 10]. WS and #Pixels denote the window size of self-attention and the corresponding number of pixels. $X$ refers to the input tensor.

### v2

@@ Abstract
In this paper, we tackle the high computational overhead of Transformers for efficient image super-resolution (SR). Motivated by the observations of self-attention’s inter-layer repetition, we introduce a convolutionized self-attention module named Convolutional Attention (ConvAttn) that emulates self-attention’s long-range modeling capability and instance-dependent weighting with a single shared large kernel and dynamic kernels. By utilizing the ConvAttn module, we significantly reduce the reliance on self-attention and its involved memory-bound operations while maintaining the representational capability of Transformers. Furthermore, we overcome the challenge of integrating flash attention into the lightweight SR regime, effectively mitigating self-attention’s inherent memory bottleneck. We scale up the window size to 32$\times$32 with flash attention rather than proposing an intricate self-attention module, significantly improving PSNR by 0.31dB on Urban100$\times$2 while reducing latency and memory usage by 16$\times$ and 12.2$\times$. Building on these approaches, our proposed network, termed Emulating Self-attention with Convolution (ESC), notably improves PSNR by 0.27 dB on Urban100$\times$4 compared to HiT-SRF, reducing the latency and memory usage by 3.7$\times$ and 6.2$\times$, respectively. Extensive experiments demonstrate that our ESC maintains the ability for long-range modeling, data scalability, and the representational power of Transformers despite most self-attention being replaced by the ConvAttn module.
11footnotetext: Corresponding author.
@@ 1 Introduction
Therefore, Transformers have gained significant attention in the SR task since they have achieved superior performance over Convolutional Neural Networks (CNNs) while requiring lower FLOPs and fewer parameters. By capturing long-range dependencies and performing input-dependent weighting through self-attention, Transformers exhibit high representational capacity with enhanced performance, especially as the training data volume increases [15, 5]. Accordingly, numerous transformer-based methods have been proposed [75, 56, 47] to capitalize on the advantages of self-attention while reducing FLOPs and parameter counts.
Figure 1: Comparison of performance, latency, and memory usage. Our methods are evaluated against representative SR models, including CNNs, Transformers, and SSMs.
@@
In this paper, we aim to design an efficiency-driven Transformer tailored for lightweight SR tasks, achieving both improved performance and memory overhead. We start by conducting a preliminary analysis of self-attention, observing that the similarity modeling performed by self-attention and resulting extracted features remain highly consistent across several layers, as shown in Figure 2. This finding exhibits that self-attention may extract overlapping features, suggesting the possibility of reducing computational overhead without compromising representational capability by utilizing efficient alternatives. Building on this finding, we establish a design strategy that retains self-attention only in the first layer of each block while replacing the remaining layers with our proposed efficient alternative, the Convolutional Attention (ConvAttn) module. To effectively emulate self-attention’s long-range modeling and instance-dependent weighting, the ConvAttn module operates with a twofold mechanism. First, it simplifies self-attention’s long-range interaction by applying convolution with a shared 13$\times$13 large kernel throughout the entire network, targeting only a subset of channels. Second, dynamic kernels are generated to capture input-dependent weighting, mimicking the adaptive nature of self-attention. By combining these components, the ConvAttn module significantly reduces the reliance on memory-intensive self-attention while maintaining the representational capability of Transformers.
With most self-attention layers replaced by ConvAttn, we leverage this efficiency to further enhance the remaining self-attention layers. Specifically, we enlarge the window size of self-attention, significantly improving performance with only a slight increase in FLOPs. However, increasing the window size leads to an enlarged score matrix, substantially raising peak memory usage. To address this, we introduce Flash Attention [10, 14] into lightweight SR tasks to avoid materializing the score matrix. Our optimized implementation allows us to scale up the window size to 32$\times$32 while reducing latency and memory usage by 16$\times$ and 12.2$\times$, respectively, as illustrated in Table 1.
@@
  • We mitigate the memory overhead of self-attention by integrating Flash Attention, thereby enlarging the window size to 32$\times$32 without excessive memory usage. To the best of our knowledge, this is the first successful application of Flash Attention in lightweight SR tasks.
@@ 2 Related Work
Traditional CNNs Early deep learning‐based SR research [13, 29, 37, 78, 77] primarily employed convolution that extracts local features from images. These approaches typically relied on stacked convolutions, which often resulted in over‐parameterization. To address this issue, numerous studies [28, 53, 35, 46] have shared their weights across multiple layers to reduce parameter size. However, many of them suffer from performance degradation since the input-independent weights are shared across multiple layers. In this paper, we share a large kernel through the entire network while leveraging input-dependently generated dynamic Depth-Wise Convolution (DWC) [22] by each layer in parallel. This approach efficiently mitigates the parameter growth of the large kernel while maintaining each layer’s representational power.
@@
Transformers Pioneering studies that significantly reduce computational costs by computing self-attention within local windows [36] or channel-wise [69] have sparked considerable interest in applying Transformers [55, 15] to the SR task. Subsequent studies have proposed methods that enhance self‐attention to strengthen its long‐range modeling capabilities while reducing both FLOPs and parameters [75, 56, 7, 82, 73, 76, 47]. Nonetheless, Transformers remain challenging to deploy due to the excessive memory access caused by self-attention and involved memory-bound operations. In this paper, we address this limitation by replacing self‐attention with convolutions that emulate its advantages, based on the observations of self-attention’s inter-layer repetition. Our approach preserves the advantages of Transformers while reducing reliance on self-attention and associated memory-bound operations.
@@
Large Kernel CNNs and State-Space Models Several recent studies have highlighted the computational burden of Transformers and investigated alternative core operators to replicate their advantages. Specifically, CNN-based models [52, 81, 63, 62] attempt to mimic the advantages of Transformers by leveraging large kernels and pixel attention [79], while State-Space Models (SSM) [20, 21] adapt their verified long-range modeling abilities from the 1D sequence domain [18, 17] to the 2D image domain. However, these approaches often overlook either the exceptional representational capability brought by self-attention’s sophisticated similarity modeling or the additional memory access caused by supplementary mechanisms to adapting 2D domains like multi-directional or attentive scanning. In this paper, we fully leverage the advantages of both self-attention and convolution while utilizing Flash Attention [10, 14] to address the memory access of self-attention without relying on complex additional mechanisms.
@@
Figure 2: Visualized inter-layer similarities of the feature extracted by self-attention ($F$) before being added to the skip path and corresponding attention map ($A^{avg}$) from SwinIR-light [36]. Both CKA [31] and cosine similarities are measured on the Urban100$\times$4 dataset. The window and head dimensions of the attention map are averaged for visualization.
@@ 3.1 Preliminary analysis
Transformers often achieve improved performance by utilizing self-attention’s long-range modeling ability and instance-dependent weighting. However, each self-attention incurs excessive memory overhead, resulting in significant latency and memory usage. Interestingly, our preliminary analysis demonstrates that the sophisticated similarity modeling of self-attention ($A^{avg}$) and the resulting extracted features ($F$) exhibit a high degree of inter-layer similarities, with averages of 89% and 87%. This observation not only aligns with recent research [80] but also explains the success of studies that reduced computational cost by sharing the attention map across layers [75, 47]. Inspired by this observation, we investigate improving efficiency without compromising representational power.
@@ 3.2 Overall Structure
Our network is composed of four main components, as shown in Figure 3. First, the LR input image $I_{LR}\in\mathbb{R}^{H\times W\times 3}$ is transformed into a shallow feature $F_{0}\in\mathbb{R}^{H\times W\times C}$ through a 3$\times$3 convolution ($\mathrm{Conv}_{3\times 3}$). Here, $H$ and $W$ denote height and width of $I_{LR}$, and $C$ denotes feature dimension. Next, $F_{0}$ and a large kernel $LK\in\mathbb{R}^{13\times 13\times 16\times 16}$ that is learned end-to-end manner and shared across the entire model are fed into the deep feature extractor $H$, and the resulting deep feature is added back to $F_{0}$ to produce the final feature $F\in\mathbb{R}^{H\times W\times C}$. In parallel, the image-wise skip module $S$ takes $I_{LR}$ as input to generate a skipped feature $F_{s}$, and the upsampler $U$ utilize both $F$ and $F_{s}$ as inputs to return the upsampled image $I_{SR}\in\mathbb{R}^{rH\times rW\times 3}$ where $r$ denotes upscaling factor. The entire process is formulated as follows:
@@ 3.3 ESC Block
Firstly, the $F_{i-1}$ is converted to the $F^{in}_{i}$ by layer normalization ($\mathrm{LN}$) and convolutional FFN ($\mathrm{ConvFFN}$[82] with the kernel size of three. Then, $\mathrm{LN}$ and naive window self-attention [36] with the window size of 32$\times$32 are employed in a residual connection to extract $F_{i,0}$. Despite its large window size, we effectively reduce memory overhead by utilizing Flash Attention. Furthermore, since $\mathrm{ConvFFN}$ is placed before, the self-attention extracts features considering local information without the complicated $QKV$ projection module. After establishing $F_{i,0}$, the block iterates residual paths composed of $\mathrm{ConvFFN}$ and proposed ConvAttn module ($\mathrm{ConvAttn}$) $M$ times. Although our methods do not employ shifted-window self-attention, these convolution-based modules after the self-attention will extract inter-window features. Finally, the result of the last iteration $F_{i,M}$ go through $\mathrm{LN}$ and $\mathrm{Conv}_{3\times 3}$, then is added to $F_{i-1}$, resulting in $F_{i}$.
@@
Convolutional Attention Module The $\mathrm{ConvAttn}$ is designed to emulate self-attention’s two key advantages: capturing long-range dependencies and instance-dependent weighting. For this purpose, $\mathrm{ConvAttn}$ leverages two types of convolution formulated as:
@@
$F^{res}_{i,j}$ $=(F^{att}_{i,j}\circledast DK_{i,j})+(F^{att}_{i,j}\circledast LK),$
$F^{fuse}_{i,j}$ $=\mathrm{Conv}^{fuse}_{1\times 1}\bigl{(}\mathrm{Concat}(F^{res}_{i,j},F^{idt}_{i,j})\bigr{)},$
where $F^{\mathrm{CF}}_{i,j}$ denotes the features extracted by $\mathrm{ConvFFN}_{i,j}$ before. We first split the $F^{\mathrm{CF}}_{i,j}$ channel-wise, resulting in $F^{att}_{i,j}\in\mathbb{R}^{H\times W\times 16}$, which consists of the 16 channels in front, and $F^{idt}_{i,j}\in\mathbb{R}^{H\times W\times C-16}$, which contains the remaining part. Two types of convolutions are only applied within $F^{att}_{i,j}$ to reduce memory access cost, unlike the methods that leverage large kernels considering only computational complexity [52, 63, 62]. Next, dynamic depth-wise kernel $DK_{i,j}\in\mathbb{R}^{3\times 3\times 1\times 16}$ [22] is created by two layers of $\mathrm{Conv}_{1\times 1}$ with global average pooling ($\mathrm{GAP}$) and gaussian error linear unit activation ($\phi$). Since enlarging $DK$’s kernel size increases the parameter size multiplied by the hidden dimension ($h$) of the kernel estimator, we fix its kernel size to 3$\times$3. Then both $DK_{i,j}$ and $LK$ are convolved on $F^{att}_{i,j}$ and sum up those, resulting convolved feature $F^{res}_{i,j}$. Based on our preliminary analysis, we hypothesize that long-range interactions do not significantly vary across layers. Therefore, we share $LK$ across the entire $\mathrm{ESCBlock}$s and $\mathrm{ConvAttn}$s, mitigating parameter growth and optimization difficulty [38, 6]. Finally, $F^{res}_{i,j}$ and $F^{idt}_{i,j}$ are concatenated channel-wise, followed by the last convolution $\mathrm{Conv}^{fuse}_{1\times 1}$ to fuse them, resulting in fused feature $F^{fuse}_{i,j}$. In this way, both the local features extracted by $\mathrm{ConvFFN}$ and the long-range features extracted by $\mathrm{ConvAttn}$ can be leveraged to capture multi-scale information.
While $DK_{i,j}$ can be merged into $LK$ by padding both spatial and channel dimensions to reduce FLOPs, we do not because it slows down latency. On the other hand, in situations where reducing FLOPs and parameter size is crucial rather than reducing latency, $LK$ can be decomposed depth-wise separable manner [25] into point-wise large kernel filter $LK^{\mathrm{c}}\in\mathbb{R}^{1\times 1\times 16\times 16}$ and depth-wise large kernel filter $LK^{\mathrm{s}}\in\mathbb{R}^{13\times 13\times 1\times 16}$. This decomposition redefines line 3 in Equation 4 as follows:
@@
In summary, $\mathrm{ConvAttn}$ efficiently emulates self-attention’s two advantages by leveraging a shared $LK$ and layer-specific $DK$s. In contrast to Transformers that share attention maps across several layers [75, 47, 80], which still require a full materialized attention map for every layer, our module only leverages a single shared $LK$, significantly reducing memory footprints. Moreover, unlike recent CNNs [52, 81, 63, 62] that use large kernel convolutions and optionally employ pixel-wise attention to inject input dependency, we employ layer-specific $DK$s, emulating self-attention’s input-dependent weight more exactly in a memory-efficient manner.
Table 1: Comparison of latency and memory usage between naive self-attention and our Flash Attention (FA) [10, 14]. WS and #Pixels denote the window size of self-attention and the corresponding number of pixels. $X$ refers to the input tensor.